### v2

@@ 5.2 Impact from Various Design Decisions
We also compare the difference between direct retrieval (DR) and solving tasks using graph queries or general-purpose languages. A deeper analysis revealed that each approach exhibits distinct strengths and weaknesses, as evidenced by the complementary performance of backend fusions. Our analysis of failed tasks indicates that, in many cases, the knowledge graph contains the required data, but the graph query fails to extract it. In such scenarios, direct retrieval, where the entire graph is included in the model’s context, performs significantly better. This is because it bypasses query composition issues. However, direct retrieval demonstrates lower accuracy in cases requiring structured, multi-step reasoning.
We also explored different tool sets, with selected results presented in Table 1. Initially, we examined the limitations of our original tools and subsequently integrated the complete Hugging Face Agents tool set into the KGoT framework, which led to improvements in accuracy, runtime, and cost efficiency. A detailed analysis allowed us to merge the most effective components from both tool sets into an optimized hybrid tool set, further enhancing accuracy and runtime while only moderately increasing costs. Key improvements include a tighter integration between the ExtractZip tool and the Text Inspector tool, which now supports Markdown, as well as enhancements to the Surfer Agent, incorporating a Wikipedia tool and augmenting viewpoint segmentation with full-page summarization. This optimized tool set was used for all subsequent experiments.
@@
We also analyzed the advantages of KGoT on different open models, see Figure 5. KGoT offers consistent advantages over HF Agents for nearly all considered models Guo et al. (2025). Interestingly, certain sizes of DeepSeek-R1 offer high zero-shot performance that outperforms both KGoT and HF Agents, illustrating potential for further improvements specifically aimed at Reasoning Language Models (RLMs) Besta et al. (2025a; b).
Figure 5: Performance on the GAIA validation set with KGoT (non-fusion) using various LLM models. For KGoT, we use Cypher graph queries for knowledge extraction from the Neo4j graph database.
Finally, we investigate the impact on performance coming from harnessing KGs, vs. using no KGs at all (the “no KG” baseline), see Figure 6. Harnessing KGs has clear advantages, with up to nearly 2$\times$ increase in the count of solved tasks. This confirms the positive impact from structuring the task related knowledge into a graph format.
6 Related Work
Our work is related to numerous LLM domains.
First, we use LangChain LangChain Inc. (2024) to facilitate the integration of the LLM agents with the rest of the KGoT system. Other such LLM integration frameworks, such as MiniChain Rush (2023) or AutoChain Forethought (2023), could be used instead.
Agent collaboration frameworks are systems such as MetaGPT Hong et al. (2024), AutoAgents Chen et al. (2024), and numerous others Zhuge et al. (2024); Tang et al. (2024); Liu et al. (2024b); Li et al. (2024); Chu et al. (2024); Wu et al. (2024); Shinn et al. (2023); Zhu et al. (2024b); Kagaya et al. (2024); Zhao et al. (2024a); Stengel-Eskin et al. (2024); Significant Gravitas (2025); Zhu et al. (2024a). The core KGoT idea can be applied to enhance such frameworks in that a KG can also be used as a common shared task representation for multiple agents solving a task together. Such a graph would be then updated by more than a single agent. Note that KGoT outperforms the highly competitive HF Agents baseline in the GAIA validation set, which means it offers more effective agent reasoning than other frameworks.
Many works exist in the domain of general prompt engineering Beurer-Kellner et al. (2024); Besta et al. (2025b); Yao et al. (2023a); Besta et al. (2024a); Wei et al. (2022); Yao et al. (2023b); Chen et al. (2023); Creswell et al. (2023); Wang et al. (2023a); Hu et al. (2024); Dua et al. (2022); Jung et al. (2022); Ye et al. (2023). One could use such schemes to further enhance respective parts of the KGoT workflow. While we already use prompts that are suited for encoding knowledge graphs, possibly harnessing other ideas from that domain could bring further benefits.
Task decomposition & planning increases the effectiveness of LLMs by dividing a task into subtasks. Examples include ADaPT Prasad et al. (2024), ANPL Huang et al. (2023), and others Zhu et al. (2024a); Shen et al. (2023). Overall, the whole KGoT workflow already harnesses recursive task decomposition: the input task is divided into numerous steps, and many of these steps are further decomposed into sub steps by the LLM agent if necessary. For example, when solving a task based on the already constructed KG, the LLM agent may decide to decompose this step similarly to ADaPT. Other decomposition schemes could also be tried, we leave this as future work.
Retrieval-Augmented Generation (RAG) is an important part of the LLM ecosystem, with numerous designs being proposed Edge et al. (2024); Gao et al. (2024); Besta et al. (2024c); Zhao et al. (2024b); Hu & Lu (2024); Huang & Huang (2024); Yu et al. (2024a); Mialon et al. (2023); Li et al. (2022); Abdallah & Jatowt (2024); Delile et al. (2024); Manathunga & Illangasekara (2023); Zeng et al. (2024); Wewer et al. (2021); Xu et al. (2024); Sarthi et al. (2024); Asai et al. (2024); Yu et al. (2024b). RAG has been used primarily to ensure data privacy and to reduce hallucinations. Using RAG is an orthogonal design choice; it could be combined with KGoT for further benefits.
Graph-Enhanced Agent Collaboration Frameworks There are works using graphs for more effective collaboration. Examples are GPTSwarm Zhuge et al. (2024), MacNet Qian et al. (2025), and AgentPrune Zhang et al. (2024). These systems differ from KGoT in that they use a graph to model and manage multiple agents in a structured way, forming a hierarchy of tools. Contrarily, KGoT uses knowledge graphs to represent the task itself, including its intermediate state. These two design choices are orthogonal and could be combined together. Moreover, while many of these systems require training, KGoT only relies on in-context learning.
Another increasingly important part of the LLM ecosystem is the usage of tools to augment the abilities of LLMs Beurer-Kellner et al. (2023); Schick et al. (2023); Xie et al. (2024). For example, ToolNet Liu et al. (2024a) uses a directed graph to model the application of multiple tools while solving a task, however focuses specifically on the iterative usage of tools at scale. KGoT harnesses a flexible and adaptable hierarchy of various tools, which can easily be extended, to solve a wide range of complex tasks.
7 Conclusion
In this paper, we introduce the Knowledge Graph of Thoughts (KGoT), an AI assistant architecture that enhances the reasoning capabilities of low-cost models while significantly reducing operational expenses. By dynamically constructing and evolving knowledge graphs (KGs) that encode the task and its resolution state, KGoT enables structured knowledge representation and retrieval, improving task success rates on complex benchmarks such as GAIA. Our extensive evaluation demonstrates that KGoT outperforms existing LLM-based agent solutions, achieving a substantial increase in task-solving efficiency of 29% or more over the competitive Hugging Face Agents baseline, while ensuring very low operational costs.
Beyond its current implementation, KGoT provides a flexible and scalable framework for AI assistant development, with potential applications in diverse fields such as automated research, data-driven decision-making, and multimodal reasoning. Future work can explore integrating KGoT with additional external tools or with advanced graph predictive schemes for more robust KG construction Besta et al. (2023a; 2024e), incorporating other classes of graph store backends such as neural graph databases Besta et al. (2022) scaling KGoT to distributed-memory clusters Blach et al. (2024), or refining its reasoning strategies by adapting more advanced task decomposition schemes.
Acknowledgements
We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines, and for their excellent technical support. We thank Timo Schneider for help with infrastructure at SPCL. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project was supported by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei Technologies. This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017103.

### v6

@@ D.2 Impact from Various Design Decisions
We explored different tool sets, with selected results presented in Table 5. Initially, we examined the limitations of our original tools and subsequently integrated the complete Hugging Face Agents tool set into the KGoT framework, which led to improvements in accuracy, runtime, and cost efficiency. A detailed analysis allowed us to merge the most effective components from both tool sets into an optimized hybrid tool set, further enhancing accuracy and runtime while only moderately increasing costs. Key improvements include a tighter integration between the ExtractZip tool and the Text Inspector tool, which now supports Markdown, as well as enhancements to the Surfer Agent, incorporating a Wikipedia tool and augmenting viewpoint segmentation with full-page summarization. This optimized tool set was used for all subsequent experiments.
@@
Figure 13: Comparison of different fusion types in respect to the task solve operation as well as the graph backend type. We report results for answering 165 GAIA validation questions across different comparison targets. DR stands for Direct Retrieval. Model: GPT-4o mini.
Graph Backend vs. Task Solve Operation We provide more detailed results in Figure 13, studying the performance of the following configurations: NetworkX + Neo4j (with query only) and NetworkX + Neo4j (with DR only) as well as Neo4j (query + DR) and NetworkX (query + DR). Overall, the fusion of backends (with DR only) offers smaller advantages than other types of fusion. This indicates that different graph querying languages have different strengths and their fusion comes with the largest combined advantage.
D.3 Runtime
We provide a runtime overview of running KGoT on the validation set of the GAIA benchmark with GPT4o-mini, Neo4j and query-based retrieval in Figure 14. The right part follows the categorization in Appendix C. We provide a more detailed analysis of the runtime in Figure 17.
Figure 14: Different runtime categorizations of the same data. Graph storage: Neo4j. Retrieval type: query. Model: GPT-4o mini.
D.4 Compute Resources
Because of the long runtime, we executed most experiments using the OpenAI API as an external resource on server compute nodes containing a AMD EPYC 7742 CPU with 128 cores running at 2.25GHz, with a total memory of 256GB. However when the LLM is called as an external resource, KGoT is able to run on commodity hardware with minimal effects on runtime.
Our experiments with locally run LLMs were executed with compute nodes containing 4x NVIDIA GH200, a respective GPU memory of 96GB, and a total memory of 896GB. In these cases, the minimum hardware requirements are dictated by the resources needed to run each LLM locally.
High-performance & scalability experiments were performed on an Apple M3 Pro with 12 cores at 4.056GHz and a total memory of 18GB.
D.5 GAIA Result Visualizations
We also implemented various automatic scripts that plot various aspects once a GAIA run is finished. In the following we provide example plots for Neo4j with query retrieval.
We provide a breakdown for each level of the GAIA benchmark into the categories that KGoT’s answers for the tasks fall into in Figure 15. We measure the runtime and costs of the various components of KGoT and illustrate them in Figure 17. We also provide insights into the tool usage, starting with the number of tasks for which a specific tools is used and whether that task was successful or not (see Figure 16). A more detailed analysis into the tool selection is provided in the plots of Figures 18 and 19 as well as the number of times the tools are used in Figure 20.
We provide now a brief explanation of the more opaque function names listed in Figure 17.
  • Any function marked as not logged refers to function or tool calls that do not incur an LLM-related cost or where usage costs are logged within the tool itself.
  • WebSurfer.forward submits a query to SerpApi.
  • Define Cypher query given new information constructs a Cypher insert query based on newly gathered information.
  • Fix JSON corrects malformed or invalid JSON for services like Neo4j.
  • Define forced retrieve queries generates a Cypher retrieval query when the maximum number of iterations is reached.
  • Generate forced solution generates a solution based on the state of the knowledge graph if no viable solution has been parsed after a Cypher retrieve or if the forced retrievals fails after exhausting all iterations.
Figure 15: Number of tasks per level that succeeded or fall into a given error category. Graph storage: Neo4j. Retrieval type: query. Model: GPT-4o mini.
Figure 16: Overview over how many tasks use a given tool and whether they are successful or not. Graph storage: Neo4j. Retrieval type: query. Model: GPT-4o mini.
(a) Cost in dollar.
(b) Number of calls.
(c) Duration in seconds.
(d) Cost per token in dollar.
(e) Cost per time in dollar/s.
(f) Tokens per second.
Figure 17: Overview over the execution time as well as the cost in dollar. Graph storage: Neo4j. Retrieval type: query. Model: GPT-4o mini.
Figure 18: Analysis of the tool selection. Graph storage: Neo4j. Retrieval type: query. Model: GPT-4o mini.
Figure 19: Analysis of the tool selection. Graph storage: Neo4j. Retrieval type: query. Model: GPT-4o mini.
Figure 20: Analysis of the tool usage. Graph storage: Neo4j. Retrieval type: query. Model: GPT-4o mini.